OpenStudy (sbentleyaz):

The position function of a particle in rectilinear motion is given by s(t) = t3 - 12t2 + 45t + 4 for t ≥ 0.Find the position and acceleration of the particle at the instant the when the particle reverses direction. Include units in your answer.

satellite73 (satellite73):

take the derivative, set it equal to zero, and solve

OpenStudy (sbentleyaz):

Thank you!

OpenStudy (sbentleyaz):

so s'(t)= 3t^2-24t+45 correct?

OpenStudy (3mar):

\[\checkmark\]

OpenStudy (sbentleyaz):

so y=0 when x=3 and when x=5, what do we do with this? we need to find the position and the acceleration. does this have anything do do with concavity??

OpenStudy (3mar):

By the way, the acceleration is the second derivative of the position function! The velocity is just the first derivative of the position function! Keep that in mind

OpenStudy (3mar):

@sbentleyaz I know you may be busy, and I am busier, so please respond faster for saving my time and yours!

OpenStudy (sbentleyaz):

so we derive it again ?

OpenStudy (3mar):

https://www.desmos.com/calculator/tlb1yxhw7j Yes, that is right to get the acceleration!

OpenStudy (sbentleyaz):

so at x=3 the acceleration is -6 and at x = 5 the acceleration is 6? @3mar

OpenStudy (3mar):

Just derive without substitution with x=3!

OpenStudy (sbentleyaz):

what do you mean

OpenStudy (3mar):

what is the function of the velocity?

OpenStudy (sbentleyaz):

3x^2-24x+45

OpenStudy (3mar):

Can you derive it now! Just derive it!

OpenStudy (sbentleyaz):

so 6x-24

OpenStudy (sbentleyaz):

is that right ?

OpenStudy (sbentleyaz):

and are we ignoring x=5?

OpenStudy (3mar):

Just the second derivative represents the accel. without substit.!

OpenStudy (sbentleyaz):

but it has to be at the instant the particle reverses direction so why dont we sub in 3

OpenStudy (sbentleyaz):

i am very confused

OpenStudy (3mar):

"at the instant the when the particle reverses direction" Where is that point in your opinion?

OpenStudy (sbentleyaz):

when the graph goes negative?

OpenStudy (3mar):

No, when the position changes its direction!

OpenStudy (sbentleyaz):

so what does that look like on a graph? is that the minimum

OpenStudy (3mar):

where is the point that has no slope at the curve of the position function? the same as the point that has a local maximum or minimum?

OpenStudy (sbentleyaz):

the local minimum

OpenStudy (3mar):

and also the local maximum, because at both the direction of motion is changed

OpenStudy (3mar):

Let me know when you are back!

OpenStudy (sbentleyaz):

Sorry I am trying to figure this out

OpenStudy (sbentleyaz):

Im still a bit confused

OpenStudy (sbentleyaz):

do we just derive it twice and sub in 3? or some other value?

OpenStudy (3mar):

I am sure that we should derive it twice but I am not sure which point would we substitute with in them!!! @jagr2713 If you could help here, I would be grateful!

OpenStudy (sbentleyaz):

would it be the point on the graph that is the minimum?

OpenStudy (3mar):

I am not sure: maximum or minimum or both!

jagr2713 (jagr2713):

I don't do calculus yet bud D:

OpenStudy (3mar):

If you could help here @triciaal, I would be grateful!

OpenStudy (sbentleyaz):

I think i figured this one out!

OpenStudy (3mar):

how?

OpenStudy (sbentleyaz):

The particle will reverse directions when the velocity is equal to zero, the first derivative of this function is the velocity equation (as stated above). so we need to find what the values of x are when y=0. Once solved, we find these to be x=3 and x=5 (as i had found before). we then use these to find the position by plugging them into the original equation. s(t) = 3^3 – 12 * 3^2 + 45 * 3 + 4 = 58 s(t) = 5^3 – 12 * 5^2 + 45 * 5 + 4 = 54 we then get 58 and 54 as the positions and we derive the velocity equation (as i found above) and get a = 6 * t – 24 we subsititue 3 and 5 in and get -6 and 6

OpenStudy (sbentleyaz):

(sorry that took me a little while to type)

OpenStudy (3mar):

No problem.

OpenStudy (3mar):

"The particle will reverse directions when the velocity is equal to zero," The velocity may be zero (as a stop) and return to be greater than zero again in the same direction!

OpenStudy (sbentleyaz):

but the graph is a parabola

OpenStudy (3mar):

So what you mean?

OpenStudy (3mar):

Let me know when you are back!